The motion of the earth's crustal plates is described by a model referred to as a plate tectonic motion.measurement indicate that the Borneo island have westward plate tectonic speed of 1.3 cm per year.estimate the time it would take for this motion to carry Borneo to Singapore island.(assume a distance of 750km)
1
Expert's answer
2018-09-18T11:34:08-0400
v=1.3 cm/year=0.013 m/year
d=750 km=750000 m.
Thus, the time it would take for this motion to carry Borneo to Singapore island is
t=d/v=750000/0.013=58 millions of years.